A new multi-fiber unit is being made with the help of G. Avila from ESO. The coating of the mirrors is expected to be finished soon, and depending on when they are finished we are preparing a more detailed planning for installing, aligning and testing of things. We are currently aiming for a visit from S. Frandsen (Aarhus) and Bo Lindberg (Lens-Tech AB) at the end of April.
The ThAr lamp on the old calibration unit (in the instrument room) is now under computer control. It was found that the split fiber that leads the light from the new calibration unit (on the adapter) to the fiber-unit entrance, gave rise to interference fringes at all wavelengths. A new type of bifurcated fiber was ordered to remedy this.